A force 800-N force is applied at an angle 30.0 degree below the horizontal to a 100-kg box initially resting on the floor. How long does it take to push the box 10.0 m to the left? The static and kinetic coefficient of friction are μs =0.4 μk = 0.3 respectively.

*Please Explain Step By Step And Also Carrying Out The Significant Figures To The Very End!*

THANK YOU!
Answer #1

Normal force , N = mg + Fsin30 = 100 x 9.8 + 800sin30 = 1380 N

force in direction of motion = Fcos30 = 800cos30 = 692.82 N

static friction force = 0.4N = 0.4 x 1380 = 552 N

kinetic friction force = 0.3 x 1380 = 414 N

now accelaration of box = (692.82 - 414)/100 = 2.7882 m/s2

d = 10 m

u = 0

d = ut + 0.5at2

10 = 0 + 0.5 x 2.7882 x t2

t = 2.6783 sec